OpenStudy (anonymous):

n a 45°- 45°- 90° right triangle, the hypotenuse is 9 cm. How long are the legs? if the other two legs are equal then it be half of the hypotenuse , 4.5 9/2,9/2

Directrix (directrix):

No. The 45-45-90 theorem states that the hypotenuse is square root of 2 times a leg. If the hypotenuse is 9, then the two congruent legs are each 9/√2 =( 9√2)/2.

OpenStudy (campbell_st):

not quite right this is an right isosceles triangle using Pythagoras' theorem  9^2 = x^2 + x^2 81 = 2x^2 \[x = \sqrt{\frac{81}{2}}\]